if a and b are real numbers, i2=-1, and (a+b)+5i=9+ai,
what are the steps to know the value of b

Answer by jim_thompson5910(35256)

Equate the imaginary parts to get 5i = ai and solve for 'a' to get a = 5


Now equate the real parts to get a+b = 9 and plug in a = 5 to get 5+b = 9


From here, solve for b:


5+b = 9


b = 9-5


b = 4


So the answer is b = 4


If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim